# AudioComposer: Towards Fine-grained Audio Generation with Natural Language Descriptions

### v1

@@ Abstract
Current Text-to-audio (TTA) models mainly use coarsetext descriptions as inputs to generate audio, which hinders models from generating audio with fine-grained control of content and style. Some studies try to improve the granularity by incorporating additional frame-level conditions or control networks. However, this usually leads to complex system design and difficulties due to the requirement for reference frame-level conditions. To address these challenges, we propose AudioComposer, a novel TTA generation framework that relies solely on natural language descriptions (NLDs) to provide both content specification and style control information. To further enhance audio generative modeling, we employ flow-based diffusion transformers with the cross-attention mechanism to incorporate text descriptions effectively into audio generation processes, which can not only simultaneously consider the content and style information in the text inputs, but also accelerate generation compared to other architectures. Furthermore, we propose a novel and comprehensive automatic data simulation pipeline to construct data with fine-grained text descriptions, which significantly alleviates the problem of data scarcity in the area. Experiments demonstrate the effectiveness of our framework using solely NLDs as inputs for content specification and style control. The generation quality and controllability surpass state-of-the-art TTA models, even with a smaller model size.  [^1] Demo and code are available in https://lavendery.github.io/AudioComposer/
@@ IV-A Results on AudioCondition
In this experiment, we train AudioComposer on all datasets introduced in Section III-A and evaluate it on AudioCondition. We compare AudioComposer with mainstream generative models, including AudioLDM [8], AudioLDM2 [11], Tango [10], Tango2 [12], and MC-Diffusion [3], to assess the performance of temporal controllability, with the results provided in Table I. As Tango and AudioLDM generate audio with no need of temporal conditions, we train Tango  [^4] https://github.com/declare-lab/tango/tree/master from scratch with language control, namely Tango+LControl. Whereas, considering the lack of style control in open-source models and the cost of human resources, we only select Tango2 for the MOS evaluation.
Table I shows that our AudioComposer outperforms these baseline models across all metrics. Interestingly, AudioComposer even surpasses the ground truth in $F1_{event}$ and $MOS_{t}$. This can be attributed to the fact that the ground truth is derived from AudioSet, which contains some extraneous noise apart from the sound events. These results collectively illustrate the outstanding performance achieved by AudioComposer, even with a smaller number of parameters in AudioComposer-S.
@@ IV-C Ablation Study
Table III provides further ablation studies on AudioCondition. Firstly, we remove flow matching and use the DDPM-based diffusion method, which results in a slight performance decline compared to AudioComposer-S. This suggests that the flow-based diffusion can enhance the model’s control capability. In terms of inference efficiency, both DDPM-based diffusion and Tango need 200 steps, while our method can generate high-quality audios with only 25 inference steps. Furthermore, when comparing our AudioComposer using DDPM-based diffusion transformer in Table III with Tango+LControl using U-Net diffusion from Table I, our diffusion transformer still achieves comparable performance, particularly in $F1_{event}$.
As demonstrated in Table III, we remove the AudioCaps data and retrain the model AudioComposer-S, which also results in a performance decline. This validates that our proposed method has the ability to utilize other coarse-grained data to enhance audio generation performance, as the models can be trained solely on NLDs in a holistic manner. Furthermore, AudioComposer can generate audios without style control, and we have evaluated it on the AudioCaps test set, as presented on our demo page due to page limitation [^5] https://lavendery.github.io/AudioComposer/.

### v2

@@ Abstract
Current Text-to-audio (TTA) models mainly use coarse text descriptions as inputs to generate audio, which hinders models from generating audio with fine-grained control of content and style. Some studies try to improve the granularity by incorporating additional frame-level conditions or control networks. However, this usually leads to complex system design and difficulties due to the requirement for reference frame-level conditions. To address these challenges, we propose AudioComposer, a novel TTA generation framework that relies solely on natural language descriptions (NLDs) to provide both content specification and style control information. To further enhance audio generative modeling, we employ flow-based diffusion transformers with the cross-attention mechanism to incorporate text descriptions effectively into audio generation processes, which can not only simultaneously consider the content and style information in the text inputs, but also accelerate generation compared to other architectures. Furthermore, we propose a novel and comprehensive automatic data simulation pipeline to construct data with fine-grained text descriptions, which significantly alleviates the problem of data scarcity in the area. Experiments demonstrate the effectiveness of our framework using solely NLDs as inputs for content specification and style control. The generation quality and controllability surpass state-of-the-art TTA models, even with a smaller model size.  [^1] Demo and code are available in https://lavendery.github.io/AudioComposer/. Work performed during an internship at Tencent AI Lab. Corresponding author. This work was supported by National Natural Science Foundation of China (62306260,62076144).
@@ IV-A Results on AudioCondition
In this experiment, we train AudioComposer on all datasets introduced in Section III-A and evaluate it on AudioCondition. We compare AudioComposer with mainstream generative models, including AudioLDM [8], AudioLDM2 [11], Tango [10], Tango2 [12], and MC-Diffusion [3], to assess the performance of temporal controllability, with the results provided in Table I. As Tango and AudioLDM generate audio without the need for temporal conditions, we train Tango  [^4] https://github.com/declare-lab/tango/tree/master from scratch with language control as a more direct comparison baseline, namely Tango+LControl. Considering the lack of style control in open-source models and the cost of human resources, we only select Tango2 for the MOS evaluation.
Table I shows that our AudioComposer outperforms these baseline models across all metrics. Interestingly, AudioComposer even surpasses the ground truth in $F1_{event}$ and $MOS_{t}$. This can be attributed to the fact that the ground truth is derived from AudioSet, which contains some extraneous noise apart from specific events. These results demonstrate the outstanding performance of AudioComposer, even with fewer parameters in AudioComposer-S.
@@ IV-C Ablation Study
Table III provides further ablation studies on AudioCondition. Firstly, we remove flow matching and use the DDPM-based diffusion method, which results in a slight performance decline compared to AudioComposer-S. This suggests that the flow-based diffusion can enhance the model’s control capability. In terms of inference efficiency, both DDPM-based diffusion and Tango need 200 steps, while our method can generate high-quality audio in just 25 steps. Our method also achieves better results with the same number of steps. Furthermore, when comparing our AudioComposer using DDPM-based diffusion transformer in Table III with Tango+LControl using U-Net diffusion from Table I, our diffusion transformer still achieves comparable performance, particularly in $F1_{event}$.
As shown in Table III, we remove the AudioCaps data and retrain the model AudioComposer-S, resulting in a performance decline. This validates that our method can utilize other coarse-grained data to enhance audio generation performance, as the models can be trained solely on NLDs in a holistic manner. Furthermore, AudioComposer can generate audio without style control, and we have evaluated it on the AudioCaps test set, as presented on our demo page due to page limitation.